### v1

# Quantum circuit generation via test-time learning with large language models

Abstract
Large language models (LLMs) can generate structured artifacts, but turning them into dependable optimizers for scientific design requires closed-loop refinement under black-box evaluation. We formulate quantum circuits synthesis as a test-time optimization problem in which an LLM proposes edits to a fixed-length gate list and an external simulator scores the resulting state using the Meyer-Wallach (MW) global entanglement measure. We introduce a lightweight test-time learning recipe that (i) reuses previously high-scoring candidates as an explicit memory trace, (ii) augment prompts with score-difference feedback, and (iii) employs restart-from-best sampling to escape plateaus. In fixed 20-qubit settings, demonstrates the feasibility of the approach but with a low success rate. Then, we move on 25 qubit, adding feedback. The full strategy mitigates a pronounced plateau observed under naive querying. Beyond raw scores, we analyze the structure of the synthesized states and find that high-MW solutions often resemble stabilizer or graph-state-like constructions, while full connectivity is not guaranteed-reflecting both properties of the MW metric and the constraints induced by prompt design. Overall, our results demonstrate the promise-and the limitations- of memory- and evaluator-guided LLM optimization for quantum circuit synthesis, and underscore the continued importance of careful objective selection and human-in-the-loop problem formulation.
I Introduction
Large language models (LLMs) have become remarkably capable of producing structured artifacts, including executable code, formal proofs, molecular strings, and hardware descriptions, where correctness is not a matter of style but of [5, 21, 10, 19, 29, llm6, llm7, llm8, 9]. Yet the central tension remains: these models are, at their core, stochastic sequence generators. This yielded to the theory of naive sampling approaches like the Large Language Monkeys [3] which, left unattended, can be inventive at the price of high computational resources. A promising direction is to move beyond one-shot prompting and instead scale inference-time [14, 17, 1, 8].
At the same time, test-time learning approaches that equip models with persistent working memory, such as dynamic cheatsheets, suggest a complementary lever: not just sampling repeatedly[3], or back-propagating feedback [28], but query with accumulated experience, to guide the model solution search.
In this work, we study these ideas in a domain that makes the reliability problem painful: quantum circuit synthesis. Prior machine learning approaches with diffusion network produced solutions for up to 9-qubit [10], and for 4,6-qubit with reinforcement learning [11, 25, 13].
In this work we explore the potential of test-learning approach, avoiding completely the need to prepare a dataset for training or fine-tuning the system. By casting the circuit synthesis as a test-time optimization loop in which an LLM proposes circuit edits from a single initially random circuit. In this way, the generation process becomes a trajectory that can accumulate “skills” specific to the task, such as reusable entangling motifs and topology patterns. We represent circuits directly as gate lists, similarly to other recent job that focus on pre-trained or fine-tuning approaches [llm6, 29], and restrict available operations to an experimentally motivated discrete subset, avoiding continuous parameters that would inflate the search space. As the validation metric, we opt for the Meyer–Wallach global entanglement measure, an experimentally-friendly and permutation-invariant metric defined in terms of single-qubit reduced purities. This choice deliberately emphasizes global structure rather than local heuristics: the model is rewarded only when the entire register becomes strongly entangled.
The presence of explicit textual feedback that encodes the change in score between iterations as a reward or penalty – for example, “You obtained an improvement/loss of about ” – provides a dense learning signal, and a restart-from-best strategy showcases greater efficiency. Starting 25-qubit, we observe a pronounced difficulty and a plateau by naive querying. Incorporating feedback and restart-from-best substantially changes this picture, steering the search toward high and extremely high entanglement solutions with a high success rate.
Taken together, our results position LLMs not merely as generative text engines but as components in closed-loop synthesis systems that: (i) generate structured candidates, (ii) incorporate external evaluation signals, and (iii) improve over time by retaining and reusing successful intermediate artifacts. We argue that such test-time memory-based learning offers a practical path toward more dependable LLM-driven design in settings where training bespoke models is infeasible, datasets are scarce, and evaluation is expensive but available. The article is structured as follows. In Section II we describe the metric used for prompting and validation, together with the inference-time approach. In Section III we display the outcomes for 20 and 25 qubits, using the GPT5.1 and 5.2 models, the role of the feedback, and the restart-from-best approach. Lastly, in Section IV we draw our observation on the method, its potential and limitations, and offer a clear future path.
II method
II.1 The simplified memory approach
The approach used in this work echoes the Dynamic Cheatsheet (CD) strategy [24]. CD is a type of test-time learning in which the model is fed the history of previous outcomes and knowledge stored in the memory. A so-called curation step is carried out in this job, cleaning potential unwanted characters, like extra parentheses or ASCII characters, generated from the model, to retrieve a clean ready to be used to prepare the new quantum circuit. [11].
II.2 The evaluation metric: Meyer-Wallach Global Entanglement
In this work, we focus on maximizing the global entanglement amount of the output state generated by each circuit topology. We exploit the Meyer-Wallach (MW) global entanglement measure in our optimization task [16] for a twofold reason: it offers an experimentally accessible and compact permutation-invariant quantifier [2, 26]; it allows us to provide a simple but clear feedback to the model prompt about the general quality of the solution, which is sampled from the ensemble of random circuits. In this way, we exploit the same idea applied in Ref. [28], where standard algorithms evaluate generated molecules, evaluating metrics of interest to refine the prompt at each iteration.
The MW is defined as follows: given a pure state of an -qubit register, living in the Hilbert space , the measure is defined as
where is the reduced density operator of the th qubit. The quantity is the purity of the reduced state, whose deviation from unity reflects the degree of mixing induced by entanglement with the rest of the register.
By construction, the measure satisfies , with one corresponding to a maximally entangled state like the GHZ, or any product of local Bell-type states with non-trivial rotations. As expected, the metric goes to zero if and only if the state is fully separable. Also, it is worth highlighting that this metric does not produce a W state.
II.3 Goal and training procedure
In real hardware, the compilation of several two-qubit gates, like the SWAP, is obtained by breaking down the unitary operation into a combination of three gates [15, 4, 18]. So, from an experimentally friendly perspective, we will fix our attention only on the following set of gates: . The gate works with a continuous angle, and this would blow up the search space where the LLM must operate. To simplify our problem, we fix the angles to be .
Solution generation
All-in-all, our circuits are not constrained and so belong to the ensemble of random quantum circuits, which we call from now on .
Starting from an initial random circuit any new query is indicated with . This is a common setting of online learning [12]. We use a lightweight test-time episodic memory, when at the th step, the model is provided with the circuit , which glean the knowledge and structure acquired from the previous step only, that is
and a feedback that offers a gain/loss feedback to the model, in analogy to the reinforcement learning approach, providing in this way a minimal curation of our memory step. This is measured as the difference in Meyer-Wallach value between steps . So we have
Similar to the Dynamics Cheatsheet [24], we curate (say, clean) the model output to extract a syntactically correct string of instructions, representing our new circuit, that can be correctly read by the algorithms to create the circuit and evaluate the amount of entanglement in it.
Prompting key steps
We prompt our model to abide by two key instructions (constraints):
Improve the Meyer-Wallach entanglement of the current circuit.
Do not add or remove more gates.
From a practical perspective, we need to keep under strict control the number of used gates, needing the minimum possible. Therefore, we aim to evaluate how well our model adheres to the instructions and to avoid trivial solutions, such as adding many random CNOT gates.
Circuit representation for the LLM – We take as the initial state for every circuit. The output state is thus the one worked out by each synthesized circuit, and then evaluated by standard algorithms. To generate new circuits, we need to represent it in a format readable by the machine. The solution is quite straightforward: we represent it as a list of gates in a string format, which encode the sequence of instructions to realize it algorithmically in common software for quantum computation. In practical terms, the LLM receive this text representation of the circuit and outputs a new one. See an example of in-train inference in Fig. 1.
II.4 Models and prompting strategy
In this work, we consider only the GPT5.1 and GPT5.2 models. We begin with considering the first one, where the prompt just uses the output circuit inside the model prompt, in the hope that the model could leverage its internal knowledge to improve it further. See the upper-left box in Fig. 1. When we move to 25 qubits, the problem computational complexity becomes evident, and we need to design a more sophisticated prompting and training strategy. Firstly, likewise reinforcement learning [23], to provide a cue for a better search direction to the model, we write an extra textual feedback that explicitly rewards (“you did well”) or penalizes (“you did worse”) the model, together with the amount of global entanglement the circuit produces. Finally, to avoid local minima, we restart the training multiple times, using the best previous outcomes as the new initial circuit.
III Main Results
Throughout the search for an optimal strategy, we tested both local, small, language models, that the most advanced at our current availability. We notice that due to the characteristic of the global metric we use, the optimal performances gravitate towards both stabilizer states [20, 22] but disconnected, or close to cluster states [6, 7] with high connectivity. As the last observation, the difference between big and small models is startling; see Appendix V.2 for their performance. Therefore, we focus only on the bleeding models available at the date, i.e., GPT5.1 and 5.2. For this first round of experiment, we focus our attention on the feedback-less prompt to see how good the model is in a task with no active guidance.
III.1 GPT5.1 for 20-qubit entangling circuits synthesis
Here, we explore the application of GPT5.1 with a memory step and no feedback. We experiment using 25 to 45 gates on 20-qubit circuits. This choice shrinks the search space, but dodges a trick recurrently applied by small models, which is the application of a high number ( num gates) of random CNOTs. Ranging from 25 to 45 gates, we can see from Table 1 that the model improves over the initial random circuit, and can obtain up to perfect MW in some situations, as we can see for 35 gates. Nevertheless, the success probability of having an MW measure is . This evidence backs up our insights, demonstrating that LLMs can sample optimal circuits beyond repeated sampling approaches, also known as the Large Language Monkeys [brown2024largelanguagemonkeysscaling]. Nevertheless, to achieve greater performances, a leap is required.
III.2 GPT5.2 for 25 qubits entangling circuits synthesis with restart-from-best and feedback.
When scaling the problem to 25 qubits, the model starts plateauing at values of MW on the newly generated circuits, for each query. Mindful of this steeple, we design two key improvements for both our prompting and testing strategy. The first one is the addition of the feedback, which introduces a further piece of information inside the prompt, in the form of a reward/penalty to the model that reads
The ratio behind this idea is to push the model to hover over its solution more consistently.
The second improvement consists in running the model more times in a row, reusing the best previous outcome at any new run. The idea behind this is that the model might get stuck in some minima, and by exploiting the statistical model behavior, we replace a perturbation with a new query. The outcomes achieved are promising. Here, we study different initial circuits under different hypotheses, i.e., randomly generated, reused, or created previously during other independent tests.
This evolution seems to support two observations. As also shown in Table 1, starting from an initial circuit that approaches ML-values can enhance solutions that are close to optimal. We argue that this effect is due to the presence of a circuit topology of higher quality, one that the model is better off exploiting. Secondly, we observe that an informative prompt can steer it towards the best solution with a higher rate of success.
Our results seem to support our observations. In Table 2 we have the results obtained from 3 different experiments, out of three attempts (no failures in between). The second row in Table 2 seems to support our speculation about the threshold ; when starting from a random circuit of extremely low amount of entanglement, the model still achieves remarkable performances but plateaus at . Instead, in rows A and C, the model can obtain almost optimal outcomes even with a single query.
Below, we report the final state synthesized by the model from experiment A, namely
This state can be rewritten as
which is a low-connected stabilizer state (other instances are provided in Appendix V.3). This is a 3-qubit GHZ multiplied by 11 Bell pairs. This circuit is entirely Clifford, so the output is a stabilizer state, LC-equivalent to some graph state.
Last, the model shows that it can also sample new states containing the gates, thus generating an almost-stabilizer state – due to the angles set choice – achieving an .
As the last check, we test over a different set of angles , to explicitly account for potential non-Clifford structure during the sampling
Table 3 demonstrates that the method also deals with non-Clifford gates. From the analysis of the outcomes (B) and (D), we notice that to attain such a high value, the LLM does not use gates, discovering Clifford solutions once more. For (A) and (C) and intermediate solutions, the rotation gates are still present.
The random-hill-climbing baseline. For a fair, budget-matched comparison, each optimization is constrained to 45 runs, matching the protocol of Table 2 (three queries with 15 candidate samples per query, where the best-scoring candidate is retained and used to warm-start the next query). In addition to the LLM-driven search, we implement a classical random-edit hill-climbing baseline that applies the same gate set and local move set (random gate replacement, wire mutation, and occasional swaps) and accepts a proposed circuit only when it improves the MW measure. Starting from the same random initialization used in Table 2, the baseline begins at and, across ten independent runs under the same 45-evaluation budget, fails to exceed . When starting from a MW measure , we achieve values of at best, always for 10 different runs. This highlights that the LLM-conditioned with feedback can discover substantially higher-scoring circuits within the same evaluation budget.
IV Conclusions
In this work, we have analyzed the ability of a new approach to quantum circuit synthesis using large language models for a global entanglement metric. We demonstrate that thinking models combined with test-time inference and feedback can obtain high values of global entanglement without any post-training. Firstly, we have designed a query with and without feedback and test its efficacy for a single run, obtaining outcomes of the order of 0.99 for 20 qubits and 35 gates, when the Meyer-Wallach metric is used to evaluate them. When moving to 25 qubits, we implement a GPT5.2 model plugging a feedback inside the prompt, and running each experiment 3 times in a row, re-using the best previous outcome to introduce a degree of perturbation. This approach immediately steers the outcomes toward peaks of MW in the synthesized solution, in a single run. This demonstrates the pivotal role played by feedback and simple memory usage for circuit synthesis with LLM, paving the way to new approaches in circuit synthesis. Interestingly, we also found that some circuits behave like ”local minima”, stopping the model from sampling improvement. This is likely due to an intrinsic limitation in the model’s reasoning ability, and is likely spawned by the lack of high-quality pre-selected theoretical knowledge, which can be used to fine-tune the model to achieve a higher level of performance and solution quality.
Future investigations will focus on designing strategies for its internal degree of connectivity, and different evaluation metrics, while from an ML perspective, the implementation of cumulative memories [24] will be considered for these more sophisticated endeavors.
Original code used for the experiment and data generated during the experiments are available at: https://github.com/AdriQD/GPT_circuits
Acknowledgements.
A.M.P. and R.L.F. acknowledge support by MUR (Ministero dell’Università e della Ricerca) through the PNRR Project ICON-Q - Partenariato Esteso NQSTI - PE00000023 - Spoke 2 - CUP: J13C22000680006.References
- [1] (2025) Inference-time computations for llm reasoning and planning: a benchmark and insights. ResearchGate Preprint. External Links: Link Cited by: §I.
- [2] (2008-03) Operational interpretation for global multipartite entanglement. Phys. Rev. Lett. 100, pp. 100503. External Links: Document, Link Cited by: §II.2.
- [3] (2024) Large language monkeys: scaling inference compute with repeated sampling. Preprint. Cited by: §I, §I.
- [4] (2025) Quantum compilation process: a survey. In Euro-Par 2024: Parallel Processing Workshops, S. Caino-Lores, D. Zeinalipour, T. D. Doudali, D. E. Singh, G. E. M. Garzón, L. Sousa, D. Andrade, T. Cucinotta, D. D’Ambrosio, P. Diehl, M. F. Dolz, A. Jukan, R. Montella, M. Nardelli, M. Garcia-Gasulla, and S. Neuwirth (Eds.), Cham, pp. 100–112. Cited by: §II.3.
- [5] (2024) A survey on evaluation of large language models. ACM Transactions on Intelligent Systems. External Links: Document, Link Cited by: §I.
- [6] (2004) Efficient algorithm to recognize the local clifford equivalence of graph states. Physical Review A 70, pp. 034302. External Links: Document, Link, quant-ph/0405023 Cited by: §III.
- [7] (2005) Local unitary versus local clifford equivalence of stabilizer states. Physical Review A 71, pp. 062323. External Links: Document, Link, quant-ph/0411115 Cited by: §III.
- [8] (2025) Inference-time computation techniques: overview and applications. Note: https://www.emergentmind.com/topics/inference-time-computationAccessed January 2026 Cited by: §I.
- [9] (2022-06) Language models can learn complex molecular distributions. Nature Communications 13 (1). External Links: ISSN 2041-1723, Link, Document Cited by: §I.
- [10] (2023) Quantum circuit synthesis with diffusion models. Preprint. Cited by: §I, §I.
- [11] (2022-10) Reinforcement-learning generation of four-qubit entangled states. Phys. Rev. Res. 4, pp. 043056. External Links: Document, Link Cited by: §I, §II.1.
- [12] (2021) Online learning: a comprehensive survey. Neurocomputing 459, pp. 249–289. External Links: ISSN 0925-2312, Document, Link Cited by: §II.3.
- [13] (2020-10) Reinforcement learning with quantum variational circuit. Proceedings of the AAAI Conference on Artificial Intelligence and Interactive Digital Entertainment 16 (1), pp. 245–251. External Links: ISSN 2326-909X, Link, Document Cited by: §I.
- [14] (2024) Adaptive inference-time compute: llms can predict if they can do better, even mid-generation. arXiv preprint arXiv:2410.02725. External Links: Link Cited by: §I.
- [15] (2022) Quantum compiling. In Quantum Computing Environments, pp. 39–74. External Links: ISBN 9783030897468, Link, Document Cited by: §II.3.
- [16] (2002-09) Global entanglement in multiparticle systems. Journal of Mathematical Physics 43 (9), pp. 4273–4278. External Links: ISSN 0022-2488, Document, Link Cited by: §II.2.
- [17] (2025) Inference-time scaling for complex tasks: where we stand and what lies ahead. Technical report Microsoft Research. External Links: Link Cited by: §I.
- [18] (2021-08) Quantum compiling by deep reinforcement learning. Communications Physics 4 (1). External Links: ISSN 2399-3650, Link, Document Cited by: §II.3.
- [19] (2025) A comprehensive overview of large language models. ACM Computing Surveys. External Links: Document, Link Cited by: §I.
- [20] (2005-12) Stabilizer formalism for operator quantum error correction. Phys. Rev. Lett. 95, pp. 230504. External Links: Document, Link Cited by: §III.
- [21] (2024) Talking about large language models. Communications of the ACM. External Links: Document, Link Cited by: §I.
- [22] (2016-10) Colloquium: protecting quantum information against environmental noise. Rev. Mod. Phys. 88, pp. 041001. External Links: Document, Link Cited by: §III.
- [23] (2018) Reinforcement learning: an introduction. Second edition, The MIT Press. External Links: Link Cited by: §II.4.
- [24] (2025) Dynamic cheatsheet: test-time learning with adaptive memory. External Links: 2504.07952, Link Cited by: §II.1, §II.3, §IV.
- [25] (2024) Reinforcement learning to disentangle multiqubit quantum states from partial observations. External Links: 2406.07884, Link Cited by: §I.
- [26] (2008) Quantum computing and entanglement for mathematicians. In Representation Theory and Complex Analysis: Lectures given at the C.I.M.E. Summer School held in Venice, Italy June 10–17, 2004, E. C. Tarabusi, A. D’Agnolo, and M. Picardello (Eds.), pp. 345–376. External Links: ISBN 978-3-540-76892-0, Document, Link Cited by: §II.2.
- [27] (2020) HuggingFace’s transformers: state-of-the-art natural language processing. External Links: 1910.03771, Link Cited by: §V.2.
- [28] (2025-03) Optimizing generative ai by backpropagating language model feedback. Nature 639 (8055), pp. 609–616. External Links: ISSN 1476-4687, Link, Document Cited by: §I, §II.2.
- [29] (2023) A survey of large language models. arXiv preprint. External Links: Link Cited by: §I, §I.
V appendix
V.1 Prompting as an interface for controllable behaviour in large language models
Large language models (LLMs) are conditional generative systems: given an input sequence (the prompt), the model produces a continuation sampled from a learned distribution. In practical deployments, prompting therefore acts as a high-leverage interface between user intent and model behaviour. Unlike fine-tuning or architectural modifications, prompting can be iterated rapidly, audited as a textual artefact, and adapted per task without retraining.
Why prompting matters.
Prompt design shapes performance by constraining the space of plausible continuations. Seemingly minor changes in instructions, context, or output format can yield substantial differences in factuality, reasoning quality, verbosity, and tone. In particular, prompts that specify (i) the intended audience and level of technicality, (ii) the required structure (for example, bullet points, tables, or a fixed schema), and (iii) task constraints (length, permissible sources, excluded content) often improve usability and consistency. Prompting is also a mechanism for risk management: explicit constraints (for example, to avoid sensitive attributes, provide uncertainty qualifiers, or request verification steps) can reduce failure modes such as overconfident fabrication and inappropriate content.
Core elements of effective prompts.
Across diverse tasks, effective prompts tend to include three ingredients: (1) goal specification (what constitutes a correct or useful answer), (2) context provision (relevant background, definitions, data, or examples), and (3) output contracts (a declared format and acceptance criteria). Output contracts are particularly important for downstream automation, enabling deterministic parsing and evaluation. Where ambiguity is likely, prompts can instruct the model to surface assumptions or ask targeted follow-up questions rather than improvising missing details.
Prompting for robustness and evaluation.
Prompt sensitivity motivates systematic evaluation. For a given task, we recommend testing prompts over representative input distributions, measuring both central tendency and tail risks (for example, rare but severe errors). Robust prompts are those whose performance degrades gracefully under paraphrase, minor perturbations to context, or changes in user phrasing. In reporting results, it is good practice to provide the exact prompts (or prompt templates) used for each experiment, alongside decoding parameters, to support reproducibility and meaningful comparison.
Our prompt analysis
Our prompt is designed to maximize the role of the LLM as an entanglement engineer. The system prompt is explicitly suited for this role, and here we file the role context and key recommendations (the model constraints) for our model to clearly define its function, and reduce rambling. In the user prompt, we prepare a chart for the model by defining:
What the input is expected to be.
Metric under consideration.
How it must work on the circuit representation.
What I must not do.
A clear explication of the ansatz, in other words, what the list represents.
Tailoring a good prompt is non-trivial, and its quality brings out important aftermath on the output quality.
V.2 Small LM
DeepSeek is a family of large language models whose performance is strongly shaped by how they allocate computation to multi-step reasoning, which looks like a “plan-then-solve” workflow: the model forms an internal scaffold, executes a stepwise derivation, and revises when contradictions or low-confidence branches appear. This strategy usually helps prioritize coherence, error correction, and robustness under ambiguity. For the baseline, we use a more standard model with lower mathematical ability to test the effect on our problem. The local models considered are: DeepSeek-R1-Distill-Llama-8B, and Zephyr-7b-beta. Few qubits problem. Main outcomes.
During our analysis, we send query to the models several times for a relatively simple problem: a 5-qubit circuit with 10-20 gates. Models setup – Both models are tested with a temperature of 0.7, to enforce originality in the sampling of new solutions, and ensure that the model will explore a greater solution space.
This solution is quite simple because the number of qubits is small, and the number of gates guarantees that good outcomes, say above 0.8 on MW, could be found (this can be easily checked by generating a few random circuits). With this setting, by using both the no-feedback approach, searching for a circuit quality improvement. Here, we test this hypothesis considering DeepSeek-R1-Distill-Llama-8B and Zephyr-7B models, both provided by Huggingface [27]. We start with a relative problem, with 4-qubit test. The main outcomes of our interest can be summarized as follows: (Ii) The Llama model achieves, on average, a 0.3 0.2 gain, starting from a random circuit of any value, while Zephir obtains 0.2 0.15. The top outcome obtain form Llama is 0.70.94 for Zephir is 0.40.5.
We notice that these local models can already offer improvements in entanglement synthesis. The prompt structure is, by and large, the following:
Nevertheless, we observe the following serious flaws:
Both models tend to create a list of gates that exceeds the maximum number of gates allowed many times, wasting more than of the queries.
The model (re)discovers the simple workaround of adding many random , to increase the entanglement amount.
Especially the second point, it is strategically correct, but exactly what we do not want out of our model, for two reasons: (i) the main goal is to uplift the initial input strictly, and (ii) what we want for a real practical problem is to use as few resources as possible. Given these setbacks, we opt for putting local models aside and solely focus on deployed GPTs called through the API.
V.3 State analysis
Here, we provide the analysis for an intermediate state obtained by GPT5.2 with feedback prompting, which achieved a value of MW of 0.8, that is
The final state obtained is disconnected, and it can be written as
Again, this is a disconnected graph state, and, of course, it is not 1-uniform because its MW entanglement is not 1.
Now we consider a state whose global entanglement amounts to 0.66, that is From its analysis, we find that the state is a highly disconnected combination of GHZ, Bell-pairs, and single qubits, with the qubits 6 and 18 of purity . The state is
Clearly, the whole state is not a graph state, nor local-Clifford equivalent to one, because of the presence of non-Clifford gates like creates a non-stabilizer entangled pair.