### v3

# Quantum circuit generation via test-time learning with large language models

Abstract
Large language models (LLMs) can generate structured artifacts, but turning them into dependable optimizers for scientific design requires closed-loop refinement under black-box evaluation. We formulate quantum circuits synthesis as a test-time optimization problem in which an LLM proposes edits to a fixed-length gate list and an external simulator scores the resulting state using the Meyer-Wallach (MW) global entanglement measure. We introduce a lightweight test-time learning recipe that (i) reuses previously high-scoring candidates as an explicit memory trace, (ii) augments prompts with score-difference feedback, and (iii) employs restart-from-best sampling to escape plateaus. In fixed 20-qubit settings, it demonstrates the feasibility of the approach, but with a low success rate. Then, we move on to 25 qubits, adding feedback. The full strategy mitigates a pronounced plateau observed under naive querying. Beyond raw scores, we analyze the structure of the synthesized states and find that high-MW solutions often resemble stabilizer or graph-state-like constructions, while full connectivity is not guaranteed, reflecting both properties of the MW metric and the constraints induced by prompt design. Overall, our results demonstrate the promise-and the limitations- of memory- and evaluator-guided LLM optimization for quantum circuit synthesis, and underscore the continued importance of careful objective selection and human-in-the-loop problem formulation.
I Introduction
Large language models (LLMs) have become remarkably capable of producing structured artifacts, including executable code, formal proofs, molecular strings, and hardware descriptions, where correctness is not a matter of style but of [4, 23, 9, 20, 32, 12, 19, 28, 8]. Yet the central tension remains: these models are, at their core, stochastic sequence generators. This yielded to the theory of naive sampling approaches like the Large Language Monkeys [brown2024monkey] which, left unattended, can be inventive at the price of high computational resources. A promising direction is to move beyond one-shot prompting and instead scale inference-time [14, 17, 21, 7].
At the same time, test-time learning approaches that equip models with persistent working memory, such as dynamic cheatsheets, suggest a complementary lever: not just sampling repeatedly[brown2024monkey], or back-propagating feedback [31], but query with accumulated experience, to guide the model solution search.
In this work, we study these ideas in a domain that makes the reliability problem painful: quantum circuit synthesis. Prior machine learning approaches with diffusion network produced solutions for up to 9-qubit [9], and for 4,6-qubit with reinforcement learning [10, 27, 13].
In this work, we explore the potential of a test-learning approach, avoiding the need to prepare a dataset for training or fine-tuning the system. By casting the circuit synthesis as a test-time optimization loop in which an LLM proposes circuit edits from a single initially random circuit. In this way, the generation process becomes a trajectory that can accumulate “skills” specific to the task, such as reusable entangling motifs and topology patterns. We represent circuits directly as gate lists, similarly to other recent jobs that focus on pre-trained or fine-tuning approaches [12, 32], and restrict available operations to an experimentally motivated discrete subset, avoiding continuous parameters that would inflate the search space. As the validation metric, we opt for the Meyer–Wallach global entanglement measure, an experimentally-friendly and permutation-invariant metric defined in terms of single-qubit reduced purities. This choice deliberately emphasizes global structure rather than local heuristics: the model is rewarded only when the entire register becomes strongly entangled.
The presence of explicit textual feedback that encodes the change in score between iterations as a reward or penalty – for example, “You obtained an improvement/loss of about $\pm\Delta$” – provides a dense learning signal, and a restart-from-best strategy showcases greater efficiency. Starting 25-qubit, we observe a pronounced difficulty and a plateau $\sim 0.7$ by naive querying. Incorporating feedback and restart-from-best substantially changes this picture, steering the search toward high and extremely high entanglement solutions with a high success rate.
Taken together, our results position LLMs not merely as generative text engines but as components in closed-loop synthesis systems that: (i) generate structured candidates, (ii) incorporate external evaluation signals, and (iii) improve over time by retaining and reusing successful intermediate artifacts. We argue that such test-time memory-based learning offers a practical path toward more dependable LLM-driven design in settings where training bespoke models is infeasible, datasets are scarce, and evaluation is expensive but available. The article is structured as follows. In Section II we describe the metric used for prompting and validation, together with the inference-time approach. In Section III we display the outcomes for 20 and 25 qubits, using the GPT5.1 and 5.2 models, the role of the feedback, and the restart-from-best approach. Lastly, in Section IV we draw our observation on the method, its potential and limitations, and offer a clear future path.
II method
II.1 The simplified memory approach
The approach used in this work echoes the Dynamic Cheatsheet (CD) strategy [26]. CD is a type of test-time learning in which the model is fed the history of previous outcomes and knowledge stored in the memory. A so-called curation step is carried out in this job, cleaning potential unwanted characters, like extra parentheses or ASCII characters, generated from the model, to retrieve a clean ready to be used to prepare the new quantum circuit. [10].
II.2 The evaluation metric: Meyer-Wallach Global Entanglement
In this work, we focus on maximizing the global entanglement amount of the output state generated by each circuit topology. We exploit the Meyer-Wallach (MW) global entanglement measure in our optimization task [16] for a twofold reason: it offers an experimentally accessible and compact permutation-invariant quantifier [1, 29]; it allows us to provide a simple but clear feedback to the model prompt about the general quality of the solution, which is sampled from the ensemble of random circuits. In this way, we exploit the same idea applied in Ref. [31], where standard algorithms evaluate generated molecules, evaluating metrics of interest to refine the prompt at each iteration.
[…]
$Q(\psi)=\frac{4}{N}\sum_{i=1}^{N}\left(1-\mathrm{Tr}\!\left[\rho_{i}^{2}\right]\right),$ (1)
[…]
By construction, the measure satisfies $0\leq Q(\psi)\leq 1$, with one corresponding to a maximally entangled state like the GHZ, or any product of local Bell-type states with non-trivial rotations. As expected, the metric goes to zero if and only if the state is fully separable. Also, it is worth highlighting that this metric does not produce a W state.
II.3 Goal and training procedure
In real hardware, the compilation of several two-qubit gates, like the SWAP, is obtained by breaking down the unitary operation into a combination of three $\mathrm{CNOT}$ gates [15, 3, 18]. So, from an experimentally friendly perspective, we will fix our attention only on the following set of gates: $\{\mathrm{CNOT,H,RY}\}$. The $\mathrm{RY}$ gate works with a continuous angle, and this would blow up the search space where the LLM must operate. To simplify our problem, we fix the angles to be $\{3,7,25\}$.
Solution generation
All-in-all, our circuits are not constrained and so belong to the ensemble of random quantum circuits, which we call from now on $\mathcal{D}_{\rm RQC}$.
Starting from an initial random circuit $\mathcal{U(\bar{\theta})}_{\rm in}\in\mathcal{D}_{\rm RQC}$ any new query is indicated with $\mathcal{U(\bar{\theta})}_{i}\sim\mathcal{D}_{\rm RQC}$. This is a common setting of online learning [11]. We use a lightweight test-time episodic memory, when at the $i$-th step, the model is provided with the circuit $\mathcal{U(\bar{\theta})}_{i}$, which glean the knowledge and structure acquired from the previous step only, that is
$\mathcal{U(\bar{\theta})}_{i+1}={\rm gen}(\mathcal{U}_{i}),$ (2)
and a feedback that offers a gain/loss feedback to the model, in analogy to the reinforcement learning approach, providing in this way a minimal curation of our memory step. This is measured as the difference in Meyer-Wallach value between steps (i,i1)(i,i-1). So we have
$\mathcal{U(\bar{\theta})}_{i+1}={\rm gen}(\mathcal{U(\bar{\theta})}_{i},\Delta Q(\psi)).$ (3)
Similar to the Dynamics Cheatsheet [26], we curate (say, clean) the model output to extract a syntactically correct string of instructions, representing our new circuit, that can be correctly read by the algorithms to create the circuit and evaluate the amount of entanglement in it.
Prompting key steps
We prompt our model to abide by two key instructions (constraints):
  • Improve the Meyer-Wallach entanglement of the current circuit.
  • Do not add or remove more gates.
From a practical perspective, we need to keep under strict control the number of used gates, needing the minimum possible. Therefore, we aim to evaluate how well our model adheres to the instructions and to avoid trivial solutions, such as adding many random CNOT gates.
[…]
Figure 1: Box (A). The training loop for the model with a prompt that does not consider the evaluation of the circuit. Box (B). The training loop with the reward. Inside the feedback, the reward in the form of a sentence you improved by this amount or a critique you made worse of this amount is provided together with the new circuit. Lower-row red boxes. An example of entangling circuit generation for 6 steps with in-train inference. On the left, the initial input circuit generated by randomly sampling from the set $\{\mathrm{CNOT,H,RY}\}$ and angle gates in $\{3,10,25\}$. In this example, the model GPT5.1 without feedback is used, with the output circuit incorporated into the prompt at each step. As explicitly stated by the prompt, the model replaces the gates with new ones in a creative way to find an original solution, i.e., avoid shortcuts like re-discovering a long list of CNOT gates, or getting stuck in a similar pattern. The final pair in the reddish right box is a product of Bell entangled pairs with non-trivial amplitude due to the $\mathrm{RY}$ rotation and some single qubit.
II.4 Models and prompting strategy
In this work, we consider only the GPT5.1 and GPT5.2 models. We begin with considering the first one, where the prompt just uses the output circuit inside the model prompt, in the hope that the model could leverage its internal knowledge to improve it further. See the upper-left box in Fig. 1. When we move to 25 qubits, the problem computational complexity becomes evident, and we need to design a more sophisticated prompting and training strategy. Firstly, likewise reinforcement learning [25], to provide a cue for a better search direction to the model, we write an extra textual feedback that explicitly rewards (“you did well”) or penalizes (“you did worse”) the model, together with the amount of global entanglement the circuit produces. Finally, to avoid local minima, we restart the training multiple times, using the best previous outcomes as the new initial circuit.
III Main Results
Throughout the search for an optimal strategy, we tested both local, small, language models, that the most advanced at our current availability. We notice that due to the characteristic of the global metric we use, the optimal performances gravitate towards both stabilizer states [22, 24] but disconnected, or close to cluster states [5, 6] with high connectivity. As the last observation, the difference between big and small models is startling; see Appendix V.2 for their performance. Therefore, we focus only on the bleeding models available at the date, i.e., GPT5.1 and 5.2. For this first round of experiment, we focus our attention on the feedback-less prompt to see how good the model is in a task with no active guidance.
III.1 GPT5.1 for 20-qubit entangling circuits synthesis
Here, we explore the application of GPT5.1 with an episodic memory step and no feedback. We experiment using 25 to 45 gates on 20-qubit circuits. This choice shrinks the search space, but dodges a trick recurrently applied by small models, which is the application of a high number ($\gg$ num gates) of random CNOTs. Ranging from 25 to 45 gates, we can see from Table 1 that the model improves over the initial random circuit, and can obtain up to perfect MW in some situations, as we can see for 35 gates. Nevertheless, the success probability of having an MW measure $Q(\psi)>0.8$ is $\sim 10\%$. This evidence backs up our insights, demonstrating that LLMs can sample optimal circuits beyond repeated sampling approaches, also known as the Large Language Monkeys [2]. Nevertheless, to achieve greater performances, a leap is required.
Table 1: Table of improvements, quantified by the MW measure $Q(\psi)$, obtained with the GPT5.1 model using the prompt without feedback (see Fig. 1). Firstly, the model complies with the prompt request and does not increase the gate number. The model is bound to offer improvement in the order of 0.5, but with an evident exception. This makes us hypothesize that a simple prompt engineering is not bound to work, and a more detailed approach might be required.
III.2 GPT5.2 for 25 qubits entangling circuits synthesis with restart-from-best and feedback.
When scaling the problem to 25 qubits, the model starts plateauing at $Q(\psi)\sim 0.48$ values of MW on the newly generated circuits, for each query. Mindful of this steeple, we design two key improvements for both our prompting and testing strategy. The first one is the addition of the feedback, which introduces a further piece of information inside the prompt, in the form of a reward/penalty to the model that reads
You obtained an improvement/loss of about $\pm\Delta Q(\psi)$ in the MW measure.
The ratio behind this idea is to push the model to hover over its solution more consistently.
The second improvement consists in running the model more times in a row, reusing the best previous outcome at any new run. The idea behind this is that the model might get stuck in some minima, and by exploiting the statistical model behavior, we replace a perturbation with a new query. The outcomes achieved are promising. Here, we study different initial circuits under different hypotheses, i.e., randomly generated, reused, or created previously during other independent tests.
Table 2: Entangling circuit synthesized with GPT5.2 inserting feedback (box (B) in Fig. 1, using 3 queries of 15 steps each with best-outcome-starter. In rows A and C, the first query is run using pre-sampled circuits. In rows B and D1, we can see two random starting circuits get stuck. It is extremely interesting that the experiments D2 and D3, where we start from D1, can never improve. We hypothesize that while the method is bound to obtain improvements quite systematically, some circuit topologies are not clearly handled by the model. These 11 outcomes are obtained from 11 runs in a row, so they represent the full collection. The wording “done” represents the end of the experiment.
This evolution seems to support two observations. As also shown in Table 1, starting from an initial circuit that approaches ML-values $Q(\psi)\sim 0.5$ can enhance solutions that are close to optimal. We argue that this effect is due to the presence of a circuit topology of higher quality, one that the model is better off exploiting. Secondly, we observe that an informative prompt can steer it towards the best solution with a higher rate of success.
Our results seem to support our observations. In Table 2 we have the results obtained from 3 different experiments, out of three attempts (no failures in between). The second row in Table 2 seems to support our speculation about the threshold $\sim 0.5$; when starting from a random circuit of extremely low amount of entanglement, the model still achieves remarkable performances but plateaus at $\sim 0.77$. Instead, in rows A and C, the model can obtain almost optimal outcomes even with a single query.
Below, we report the final state synthesized by the model from experiment A, namely [(’H’, [0]), (’CNOT’, [0, 12]), (’H’, [1]), (’CNOT’, [1, 13]), (’H’, [2]), (’CNOT’, [2, 14]), (’H’, [3]), (’CNOT’, [3, 15]), (’H’, [4]), (’CNOT’, [4, 16]), (’H’, [5]), (’CNOT’, [5, 17]), (’H’, [6]), (’CNOT’, [6, 18]), (’H’, [7]), (’CNOT’, [7, 19]), (’H’, [8]), (’CNOT’, [8, 20]), (’H’, [9]), (’CNOT’, [9, 21]), (’H’, [10]), (’CNOT’, [10, 22]), (’H’, [11]), (’CNOT’, [11, 23]), (’CNOT’, [23, 24])].
This state can be rewritten as
[…]
which is a low-connected stabilizer state (other instances are provided in Appendix V.3). This is a 3-qubit GHZ multiplied by 11 Bell pairs. This circuit is entirely Clifford, so the output is a stabilizer state, LC-equivalent to some graph state.
Last, the model shows that it can also sample new states containing the $\mathrm{RY}$ gates, thus generating an almost-stabilizer state – due to the angles set choice – achieving an $Q(\psi)=0.99$. [(’H’, [0]), (’CNOT’, [0, 1]), (’RY’, [25.0, 2]), (’CNOT’, [1, 2]), (’CNOT’, [2, 3]), (’RY’, [10.0, 3]), (’H’, [4]), (’CNOT’, [4, 5]), (’CNOT’, [5, 6]), (’CNOT’, [6, 7]), (’CNOT’, [7, 8]), (’H’, [9]), (’CNOT’, [9, 10]), (’RY’, [25.0, 10]), (’H’, [11]), (’CNOT’, [10, 11]), (’CNOT’, [11, 12]), (’CNOT’, [12, 13]), (’CNOT’, [13, 14]), (’RY’, [10.0, 14]), (’CNOT’, [14, 15]), (’CNOT’, [15, 16]), (’CNOT’, [16, 17]), (’RY’, [25.0, 18]), (’CNOT’, [17, 18]), (’CNOT’, [18, 19]), (’CNOT’, [19, 20]), (’RY’, [10.0, 20]), (’CNOT’, [20, 21]), (’CNOT’, [21, 22]), (’CNOT’, [22, 23]), (’RY’, [25.0, 24]), (’CNOT’, [23, 24]), (’CNOT’, [8, 16]), (’CNOT’, [3, 21]), (’RY’, [3.0, 6])].
As the last check, we test over a different set of $\mathrm{RY}$ angles $\{0.1,0.42,1.0\}$, to explicitly account for potential non-Clifford structure during the sampling
Table 3: Entangling circuit synthesized with GPT5.2 using box (B) in Fig. 1 for the prompt and same approach used for Table 2, with the only difference that for $\mathrm{RY}$ the set of angles considered is $\{0.1,0.42,1\}$. In this way, we also have non-Clifford circuits. For our collection, we show the 4 outcomes obtained from 4 syntheses.
Table 3 demonstrates that the method also deals with non-Clifford gates. From the analysis of the outcomes (B) and (D), we notice that to attain such a high value, the LLM does not use $\mathrm{RY}$ gates, discovering Clifford solutions once more. For (A) and (C) and intermediate solutions, the rotation gates are still present.
The random-hill-climbing baseline. For a fair, budget-matched comparison, each optimization is constrained to 45 runs, matching the protocol of Table 2 (three queries with 15 candidate samples per query, where the best-scoring candidate is retained and used to warm-start the next query). In addition to the LLM-driven search, we implement a classical random-edit hill-climbing baseline that applies the same gate set and local move set (random gate replacement, wire mutation, and occasional swaps) and accepts a proposed circuit only when it improves the MW measure. Starting from the same random initialization used in Table 2, the baseline begins at $Q(\psi)\simeq 0.02$ and, across ten independent runs under the same 45-evaluation budget, fails to exceed $Q(\psi)\simeq 0.16$. When starting from a MW measure $Q(\psi)=0.08$, we achieve values of $Q(\psi)\simeq 0.29$ at best, always for 10 different runs. This highlights that the LLM-conditioned with feedback can discover substantially higher-scoring circuits within the same evaluation budget.
IV Conclusions
In this work, we have analyzed the ability of a new approach to quantum circuit synthesis using large language models for a global entanglement metric. We demonstrate that thinking models combined with test-time inference and feedback can obtain high values of global entanglement without any post-training. Firstly, we have designed a query with and without feedback and test its efficacy for a single run, obtaining outcomes of the order of 0.99 for 20 qubits and 35 gates, when the Meyer-Wallach metric is used to evaluate them. When moving to 25 qubits, we implement a GPT5.2 model plugging a feedback inside the prompt, and running each experiment 3 times in a row, re-using the best previous outcome to introduce a degree of perturbation. This approach immediately steers the outcomes toward peaks of MW in the synthesized solution, in a single run. This demonstrates the pivotal role played by feedback and a simple episodic memory usage for circuit synthesis with LLM, paving the way to new approaches in circuit synthesis. Interestingly, we also found that some circuits behave like ”local minima”, stopping the model from sampling improvement. This is likely due to an intrinsic limitation in the model’s reasoning ability, and is likely spawned by the lack of high-quality pre-selected theoretical knowledge, which can be used to fine-tune the model to achieve a higher level of performance and solution quality.
Future investigations will focus on designing strategies for its internal degree of connectivity, and different evaluation metrics, while from an ML perspective, the implementation of cumulative memories [26] will be considered for these more sophisticated endeavors.
Original code used for the experiment and data generated during the experiments are available at: https://github.com/AdriQD/GPT_circuits.
Acknowledgements.
A.M.P. and R.L.F. acknowledge support by MUR (Ministero dell’Università e della Ricerca) through the PNRR Project ICON-Q - Partenariato Esteso NQSTI - PE00000023 - Spoke 2 - CUP: J13C22000680006.